The sides of one triangle are 12 inches, 14 inches, and 15 inches long, respectively. In a second triangle similar to the first, the shortest side is 8 inches long. To the nearest tength of an inch, what is the legth of the longest side of the second triangle? a) 64, b) 9.3, c) 10.0, d) 11.0, e) 14.4
Answer by shree840(260)

12/8=14/x=15/y
or 3/2=14/x=15/y
or x=28/3 y =30/3=10
Ans 28/3 and 10